OpenStudy (anonymous):

Solve:

OpenStudy (anonymous):

\[N + \frac{ 1 }{ 2 } = \frac{ 5}{ 6 }\]

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (abmon98):

get rid of the fraction by multiplying by 6 6N+3=5 6N=5-3 6N=2/6 N=2/6

OpenStudy (anonymous):

move 1/2 to other side \[N=\frac{ 5 }{ 6 } -\frac{ 1 }{ 2 }\] then find the L..C.M of the Denominator \[N=\frac{ 5-3 }{ 6 }\] N=2/6 =1/3

OpenStudy (anonymous):

Thanks! @Abdulhameed

OpenStudy (anonymous):

YW